OpenStudy (anonymous):

solve for x x^4 + 2x^2 = 3

OpenStudy (anonymous):

Let z=x^2 Than you have z^2+2z-3=0 (z+3)(z-1)=0 so z=1 or z=-3 1=x^2 implies x=1 or x=-1 -3=x^2 has no solution